Biochem help
media%2F922%2F922084e7-67ae-400c-b9fa-c2
molecule c molecule d QUESTION 6 Absorption of drugs into the body is based on multiple factors including ionization state, surface area of the stomach and intestines, and other physical properties of the drug. A novel drug contains a carboxylic acid functional group with a pKa value of 3.24.. Based on ionization states alone, would you predict that it is better absorbed (ie passes through a lipid membrane) in the stomach where the pH is about 2, or in the small intestines where the pH is about 8? In the stomach, because the acid form is predominantly present. In the stomach because the conjugate base is predominantly present. In the small intestines, because the acid form is predominantly present In the small intestines, because the conjugate base is predominantly present 0.2 QUESTION 7 A dialysis bag contianing 2M glucose is placed in a 4.5M NaCl solution. In which direction does the water flow.
QUESTION 9 Which range of pH is buffered by a oxalatic acid/sodium oxalate buffer? 6.5-7.5 3-5 3.3-5.3 4.2-4.4 QUESTION 10 Which of the following would act as a buffer system? Select all that apply 3NH4 Cr CH3COOH and HCI

molecule a molecule b H-C H,C. H C H C molecule c molecule d Which of the following could be used to form bilayers. Select all that apply molecule a molecule b molecule c molecule d
molecule a molecule b H-C H.C H.C N н.с н.с HC mo lecule c molecule d Which of the following have polar properties? Select all that apply. molecule a molecule b molecule c molecule d
molecule a molecule b H-C St H.C H,C HC H C H,C - molecule c molecule d Which of the following have non-polar properties? Select all that apply molecule a molecule b molecule c molecule d 0.2 point
Answer #1

question 11)

pH = pKa + log ( A- / HA)

HA = NaH2PO4

A- = Na2HPO4

pH = pKa + log ( Na2HPO4 / NaH2PO4)

pKa of NaH2PO4 = 7.2

pH = 7

7 = 7.2 + log ( Na2HPO4 / NaH2PO4)

log ( Na2HPO4 / NaH2PO4) = 7 - 7.2

log ( Na2HPO4 / NaH2PO4) = - 0.2

Na2HPO4 / NaH2PO4 = 10^(-0.2)

Na2HPO4 / NaH2PO4 = 0.631

NaH2PO4 / Na2HPO4 = 1 /  0.631

NaH2PO4 / Na2HPO4 = 1.5

so the ratio is 1.5 NaH2PO4 / 1.0 Na2HPO4